# A positive density of elliptic curves are diophantine stable in certain Galois Extensions

### v1

@@ 3.1. Controlling the torsion
An elliptic curve $E_{/\mathbb{Q}}$ gives rise to a family of Galois representations. Choose an algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$ and set $\operatorname{G}_{\mathbb{Q}}:=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Given a natural number $n$, let $E[n]$ be the $n$-torsion subgroup of $E(\overline{\mathbb{Q}})$. Note that $E[n]\simeq\left(\mathbb{Z}/n\mathbb{Z}\right)^{2}$ and that there is a natural action of $\operatorname{G}_{\mathbb{Q}}$ on $E[n]$. The associated Galois representation
$\rho_{E,n}:\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{Aut}(E[n])\xrightarrow{\sim}\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})$
encodes many arithmetic properties of $E$. Suppose that $E$ does not have complex multiplication. Then, Serre’s Open image theorem implies that for all but finitely many primes $\ell$, the $-\ell$ representation $\rho_{E,\ell}$ is surjective. On the other hand, a prime $\ell$ for which $\rho_{E,\ell}$ fails to be surjective is called an exceptional prime. In this context, we set $S_{1}^{\prime}\subset\mathcal{C}$ to be the set of isomorphism classes of elliptic curves with no exceptional primes.
Theorem 3.5 (Duke [Duk97]).
@@
$\rho_{E,\ell}:\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$
@@
By way of contradiction, assume that $E(L)[\ell]\neq 0$. Thus, $\rho_{E,\ell}(\operatorname{G}_{L})$ is contained in the mirabolic subgroup $M:=\left\{\left({\begin{array}[]{cc}1&\ast\\ 0&\ast\\ \end{array}}\right)\right\}$ of $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$. In particular, we find that
$[\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z}):\rho_{E,\ell}(\operatorname{G}_{L})]\geq[\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z}):M]=\frac{(\ell^{2}-\ell)(\ell^{2}-1)}{(\ell-1)\ell}=\ell^{2}-1.$
@@
$[\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z}):\rho_{E,\ell}(\operatorname{G}_{L})]=[\rho_{E,\ell}(\operatorname{G}_{\mathbb{Q}}):\rho_{E,\ell}(\operatorname{G}_{L})]\leq[\operatorname{G}_{\mathbb{Q}}:\operatorname{G}_{L}]\leq 5.$
@@
Recall that Theorem 3.5 asserts that the density of the set $S_{1}^{\prime}$ is $1$. Hence, to prove the result, it suffices to show that $S_{1}^{\prime}\subset S_{1}$. Let $E$ be an elliptic curve in $S_{1}^{\prime}$. This means by definition, that for all primes $\ell$, the Galois representation $\rho_{E,\ell}:\operatorname{G}_{\mathbb{Q}}\to\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ is surjective. Then by Lemma 3.6, we get that $E(L)[\ell]=0$ for all primes $\ell$. This in turn implies that $E(L)_{\operatorname{tors}}=0$ and hence $E\in S_{1}$. This completes the proof. ∎
@@ 3.2. Controlling the rank
In this section, we take $S_{2}\subset\mathcal{C}$ to consist of isomorphism classes of elliptic curves $E_{/\mathbb{Q}}$ such that $\operatorname{rank}E(L)=0$. Given a pair $(\overline{A},\overline{B})\in(\mathbb{Z}/\ell\mathbb{Z})^{2}$ with $4\overline{A}^{3}+27\overline{B}^{2}\neq 0$, let $E_{\overline{A},\overline{B}}:y^{2}=x^{3}+\overline{A}x+\overline{B}$. Let $\ell$ be a prime that ramifies in $L$, then set $\mathfrak{A}_{\ell}$ to denote the number of pairs $(\overline{A},\overline{B})\in(\mathbb{Z}/\ell\mathbb{Z})^{2}$ such that
@@
    for all $\ell\in Z$, there exists an elliptic curve $\mathbb{E}$ over $\mathbb{F}_{\ell}$, such that $\mathbb{E}(\mathbb{F}_{\ell})[p]=0$ and $\mathbb{E}^{-1}(\mathbb{F}_{\ell})[p]=0$. Here, $\mathbb{E}^{-1}$ is the quadratic twist of $\mathbb{E}$ by $-1$.
@@
$\eta_{p}:=\begin{cases}\frac{1}{4}&\text{ if }p=3;\\ \frac{3}{8}&\text{ if }p=5;\end{cases}$
@@
$\delta_{\ell}:=\begin{cases}\frac{2}{3}&\text{ if }\ell\notin Z\cup\{2,p\}\text{ and }\ell=3;\\ 1-\frac{2}{\ell^{2}}+\frac{1}{\ell^{3}}&\text{ if }\ell\notin Z\cup\{2,p\}\text{ and }\ell\neq 3;\\ 1-\frac{1}{\ell}&\text{ if }\ell=p;\\ \frac{1}{2^{21}}&\text{ if }\ell=2;\\ \frac{\#\mathfrak{A}_{\ell}}{\ell^{2}}&\text{ if }\ell\in Z.\end{cases}$
We note that (2) is a condition on the prime $\ell$, and it is shown that when $p=5$ and $\ell\geq 5779$ (resp. $p=3$ and $\ell\equiv 1\pmod{4}$), this condition is satisfied, see Lemma 4.6. We shall also give explicit lower bounds for the density of the set of elliptic curves that are diophantine stable in $L$ (in terms of $E$ and the set of primes that ramify in $L$). We shall show that $\operatorname{Sel}_{p}(E/L)=0$ for a positive density of elliptic curves $E_{/\mathbb{Q}}$. This is achieved by applying Proposition 2.7 by showing that Assumption 2.1 (for $K:=\mathbb{Q}$ and $p\in\{3,5\}$) is satisfied for a positive density of elliptic curves. Recall that $p$ is unramified in $L$, and hence, Assumption 2.1 (3) is satisfied.
@@
Let $E_{A,B}:y^{2}=x^{3}+Ax+B$ be an elliptic curve such that either $\ell^{4}\nmid A$ or $\ell^{6}\nmid B$. Assume that choice of $\Sigma_{\ell}$ is made for every prime number. Then, we set $\Sigma$ to denote the tuple of conditions $(\Sigma_{\ell})_{\ell}$, where $\ell$ ranges over all prime numbers. We associate the family $\mathcal{F}_{\Sigma}$ of elliptic curves to the tuple $(\Sigma_{\ell})_{\ell}$, where $E_{A,B}\in\mathcal{F}_{\Sigma}$ if $(A,B)\in\Sigma_{\ell}$ for all $\ell$. The family $\mathcal{F}_{\Sigma}$ is then said to be defined by congruence conditions.
@@
$\operatorname{Inv}(\mathcal{F}):=\{(A,B):E_{A,B}\in\mathcal{F}\}$
@@
Recall that the root number $\omega(E)$ of an elliptic curve $E_{/\mathbb{Q}}$ is the sign of the functional equation of the $L$-function $L(E,s)$ of $E$. It is widely believed that when the elliptic curves are ordered by height, the root numbers $+1$ and $-1$ occur equally often. The following result of Bhargava and Shankar [BS15b, Theorem 41] then gives the density of elliptic curves with rank 00.
@@
Let $p$ be a prime in $\{3,5\}$. Suppose $\mathcal{F}$ is a large family of elliptic curves such that exactly $50\%$ of the curves in $\mathcal{F}$, when ordered by height, have root number $+1$. Then if $p=3$ (resp. $p=5$) at least $25\%$ (resp. $37.5\%$) of the curves $E_{A,B}$ in $\mathcal{F}$, when ordered by height, have $\operatorname{Sel}_{p}(E_{A,B}/\mathbb{Q})=0$.
@@
The statement of the result above is a little stronger than that of [BS15b, Theorem 41] (resp. [BS13, Proposition 40]) since it is asserted that $\operatorname{Sel}_{p}(E_{A,B}/\mathbb{Q})=0$. The proof of the mentioned Theorems rely on showing that this Selmer group is indeed 00 for at least $25\%$ (resp. $37.5\%$) of curves in $\mathcal{F}$. When $p=3$, we refer to p. 617, ll. 11-12 of [BS15b]. On the other hand, for $p=5$, see [BS13, p.29, l. -7]. ∎
For an elliptic curve $E:=E_{A,B}:y^{2}=x^{3}+Ax+B$ with $A,B\in\mathbb{Z}$, denote by $E^{-1}:=E^{-1}_{A,B}:y^{2}=x^{3}+Ax-B$ the twist by $-1$ of $E_{A,B}$. Note that $\Delta(E^{-1})=\Delta(E)$ and $j(E^{-1})=j(E)$.
@@
(3.2) $\omega_{\ell}(E)=\begin{cases}+1&\text{ if }E\text{ has good reduction or non-split multiplicative reduction at }\ell;\\ -1&\text{ if }E\text{ has split multiplicative reduction at }\ell.\end{cases}$
There is a simple criterion for this at a prime $\ell\geq 3$ of multiplicative reduction. The elliptic curve $E=E_{A,B}$ has split multiplicative reduction precisely when $\left(\frac{6B}{\ell}\right)=1$ (see [Won01, Lemma 10]). One has that $\left(\frac{-1}{\ell}\right)=1$ if and only if $\ell\equiv 1\pmod{4}$. Therefore, we deduce that for any prime $\ell\geq 3$ at which $E$ has multiplicative reduction,
@@
    Let $\Delta^{\prime}(E)$ be the positive prime to $2$ part of $\Delta(E)$ (i.e., $\Delta^{\prime}(E):=\frac{|\Delta(E)|}{2^{v_{2}(\Delta(E))}}$). Then, $\Delta^{\prime}(E)$ is squarefree and $\Delta^{\prime}(E)\equiv 1\pmod{4}$.
@@
Let $E$ be an elliptic curve in $\mathcal{E}^{0}$. Then, setting $K:=\mathbb{Q}$, the Assumption 2.1 is satisfied for $\mathcal{E}^{0}$.
@@
  • Finally, let us consider $\ell=2$. By part (4) of Definition 3.12, $E$ has additive reduction at $2$ and $2\nmid j(E)$. It follows from [Won01, Lemma 12] that $\omega_{2}(E^{-1})=-\omega_{2}(E)$.
@@
Recall that the height of $E$, denoted $h(E)$, is defined to be $\operatorname{max}\{|A|^{3},B^{2}\}$. Thus, $h(E)$ remains unaltered under the twist by $-1$, i.e. $h(E)=h(E^{-1})$. Hence, Lemma 3.15 implies that exactly $50\%$ of the curves $E$ in $\mathcal{E}$ have root number $+1$, when ordered by height. We note that there is a natural short exact sequence
@@ 4.1. A suitable large family of elliptic curves
Clearly, the set $\mathcal{F}$ is a large family, and at the primes $T:=Z\cup\{2,p\}$, we have associated congruence conditions.
@@
for all $E\in\mathcal{F}$. That $E$ and $E^{-1}$ have additive reduction at $2$ follows from Tate’s algorithm (see [Sil94, step 7, p. 367]). Since it is a standard procedure, we omit these details. The $j$-invariant is given as follows
@@
    Given $X>0$, we set $\mathcal{F}_{\ell}(X):=\left\{(A,B)\in\mathcal{F}_{\ell}\mid\operatorname{max}\{|A|^{3},B^{2}\}\leq X\right\}$. We note that not all pairs $(A,B)\in\mathcal{F}_{\ell}$ are minimal, but the pairs $(A,B)\in\mathcal{F}$ are. We derive an asymptotic formula for #$\#\mathcal{F}_{\ell}(X)$ as $X\rightarrow\infty$.
@@
    Next, we show that the asymptotic formulae for #$\#\mathcal{F}_{\ell}(X)$ can be used to give an asymptotic formula for #$\#\mathcal{F}(X)$. Using this formula, we derive an expression for the density of $\mathcal{F}$. Since $\mathcal{F}$ is contained in $\mathcal{E}$, this gives us lower bound for the lower density of $\mathcal{E}$.
In order to complete step (1) in an effective manner, we introduce the residue classes ¯$\overline{\mathcal{F}}_{\ell}\subset(\mathbb{Z}/\ell^{n_{\ell}})^{2}$ associated to the local condition $\mathcal{F}_{\ell}$ for all primes. We then set $\delta_{\ell}:=\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}$. The asymptotic for #$\#\mathcal{F}_{\ell}(X)$ is then given by
#$\#\mathcal{F}_{\ell}(X)\sim\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}\times\#\{(A,B)\in\mathbb{Z}^{2}\mid\operatorname{max}\{|A|^{3},B^{2}\}\leq X\}\sim 4\delta_{\ell}X^{5/6}.$
@@
  • For $\ell\notin T$, we set $n_{\ell}=2$ and ¯$\overline{\mathcal{F}}_{\ell}\subset(\mathbb{Z}/\ell^{2}\mathbb{Z})^{2}$ to consist of pairs $(\overline{A},\overline{B})$ for which $4\overline{A}^{3}+27\overline{B}^{2}\neq 0$.
@@
  • For $\ell=2$, we set $n_{\ell}=12$ and ¯$\overline{\mathcal{F}}_{\ell}\subset(\mathbb{Z}/2^{12}\mathbb{Z})^{2}$ to consist of pairs $(\overline{A},\overline{B})$ where $(\overline{A},\overline{B})=(756,\pm 16)$.
@@
Let $\ell$ be a prime, we find that #$\#\mathcal{F}_{\ell}(X)\sim 4\delta_{\ell}X^{5/6}$, where
$\delta_{\ell}=\begin{cases}\frac{2}{3}&\text{ if }\ell\notin T\text{ and }\ell=3;\\ 1-\frac{2}{\ell^{2}}+\frac{1}{\ell^{3}}&\text{ if }\ell\notin T\text{ and }\ell\neq 3;\\ 1-\frac{1}{\ell}&\text{ if }\ell=p;\\ \frac{1}{2^{21}}&\text{ if }\ell=2.\\ \end{cases}$
@@
¯$\overline{\mathcal{F}}_{\ell}\subset(\mathbb{Z}/\ell^{n_{\ell}})^{2}$
@@
#$\#\mathcal{F}_{\ell}(X)\sim\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}\right)\times 4X^{5/6}.$
We count $\#\overline{\mathcal{F}}_{\ell}$ in all cases.
  • First consider $\ell\notin T$. Recall that for $\ell\notin T$, $\mathcal{F}_{\ell}$ consists of all pairs $(A,B)\in\mathbb{Z}^{2}$ such that $\ell^{2}\nmid\Delta_{A,B}$. Thus, $\overline{\mathcal{F}}_{\ell}$ consists of all pairs $(\overline{A},\overline{B})\in(\mathbb{Z}/\ell^{2})^{2}$ such that $-4\overline{A}^{3}\neq 27\overline{B}^{2}$. Note that $\ell\neq 2$ since $2\in T$. First assume that $\ell=3$ (this case potentially arises when $p=5$). In this case, the condition becomes $3\nmid\overline{A}$, and thus, #¯$\#\overline{\mathcal{F}}_{\ell}=2\times 3^{3}$ and thus,
    (#¯$\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}\right)=\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{4}}\right)=\frac{2}{3}.$
    Next, we move on the case when $\ell\neq 3$. Set $(C,D):=(-\overline{A}/3,\overline{B}/2)$ and count the number of pairs for which $C^{3}=D^{2}$. We consider two further cases. In the first case, $\ell\nmid C$, we set $(U,V):=(C,D/U)$ and the equation becomes $U=V^{2}$. Thus, in this case, there are $(\ell^{2}-\ell)$ values of $V$ and $U$ is determined by $V$. On the other hand, if $\ell|C$ then $\ell|D$ as well and there are $\ell^{2}$ choices in this case. Thus, we get
    #¯$\#\overline{\mathcal{F}}_{\ell}=\ell^{4}-(\ell^{2}-\ell)-\ell^{2}=\ell^{4}-2\ell^{2}+\ell.$
@@
    (#¯$\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}\right)=\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{4}}\right)=1-\frac{2}{\ell^{2}}+\frac{1}{\ell^{3}}.$

### v2

@@ 3.1. Controlling the torsion
An elliptic curve $E_{/\mathbb{Q}}$ gives rise to a family of Galois representations. Choose an algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$ and set $\operatorname{G}_{\mathbb{Q}}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Given a natural number $n$, let $E[n]$ be the $n$-torsion subgroup of $E(\overline{\mathbb{Q}})$. Note that $E[n]\simeq\left(\mathbb{Z}/n\mathbb{Z}\right)^{2}$ and that there is a natural action of $\operatorname{G}_{\mathbb{Q}}$ on $E[n]$. The associated Galois representation
$\rho_{E,n}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{Aut}(E[n])\xrightarrow{\sim}\operatorname{GL}_{2}(\mathbb{Z}/n\mathbb{Z})$
encodes many arithmetic properties of $E$. Suppose that $E$ does not have complex multiplication. Then, Serre’s Open image theorem implies that for all but finitely many primes $\ell$, the mod-$\ell$ representation $\rho_{E,\ell}$ is surjective. On the other hand, a prime $\ell$ for which $\rho_{E,\ell}$ fails to be surjective is called an exceptional prime. In this context, we set $S_{1}^{\prime}\subset\mathcal{C}$ to be the set of isomorphism classes of elliptic curves with no exceptional primes.
Theorem 3.5 (Duke [dukeexceptional]).
@@
$\rho_{E,\ell}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$
@@
By way of contradiction, assume that $E(L)[\ell]\neq 0$. Thus, $\rho_{E,\ell}(\operatorname{G}_{L})$ is contained in the mirabolic subgroup $M\mathrel{\mathop{\ordinarycolon}}=\left\{\left({\begin{array}[]{cc}1&\ast\\ 0&\ast\\ \end{array}}\right)\right\}$ of $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$. In particular, we find that
$\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})\mathrel{\mathop{\ordinarycolon}}\rho_{E,\ell}(\operatorname{G}_{L})]\geq[\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})\mathrel{\mathop{\ordinarycolon}}M]=\frac{(\ell^{2}-\ell)(\ell^{2}-1)}{(\ell-1)\ell}=\ell^{2}-1.$
@@
$\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})\mathrel{\mathop{\ordinarycolon}}\rho_{E,\ell}(\operatorname{G}_{L})]=[\rho_{E,\ell}(\operatorname{G}_{\mathbb{Q}})\mathrel{\mathop{\ordinarycolon}}\rho_{E,\ell}(\operatorname{G}_{L})]\leq[\operatorname{G}_{\mathbb{Q}}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{L}]\leq 5.$
@@
Recall that Theorem 3.5 asserts that the density of the set $S_{1}^{\prime}$ is $1$. Hence, to prove the result, it suffices to show that $S_{1}^{\prime}\subset S_{1}$. Let $E$ be an elliptic curve in $S_{1}^{\prime}$. This means by definition, that for all primes $\ell$, the Galois representation $\rho_{E,\ell}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\to\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ is surjective. Then by Lemma 3.6, we get that $E(L)[\ell]=0$ for all primes $\ell$. This in turn implies that $E(L)_{\operatorname{tors}}=0$ and hence $E\in S_{1}$. This completes the proof. ∎
@@ 3.2. Controlling the rank
In this section, we take $S_{2}\subset\mathcal{C}$ to consist of isomorphism classes of elliptic curves $E_{/\mathbb{Q}}$ such that $\operatorname{rank}E(L)=0$. Given a pair $(\overline{A},\overline{B})\in(\mathbb{Z}/\ell\mathbb{Z})^{2}$ with $4\overline{A}^{3}+27\overline{B}^{2}\neq 0$, let $E_{\overline{A},\overline{B}}\mathrel{\mathop{\ordinarycolon}}y^{2}=x^{3}+\overline{A}x+\overline{B}$. Let $\ell$ be a prime that ramifies in $L$, then set $\mathfrak{A}_{\ell}$ to denote the number of pairs $(\overline{A},\overline{B})\in(\mathbb{Z}/\ell\mathbb{Z})^{2}$ such that
@@
    for all $\ell\in Z$, there exists an elliptic curve $\mathbb{E}$ over $\mathbb{F}_{\ell}$, such that $\mathbb{E}(\mathbb{F}_{\ell})[p]=0$ and $\mathbb{E}^{-1}(\mathbb{F}_{\ell})[p]=0$. Here, $\mathbb{E}^{-1}$ is the quadratic twist of $\mathbb{E}$ by 1-1.
@@
$\eta_{p}\mathrel{\mathop{\ordinarycolon}}=\begin{cases}\frac{1}{4}&\text{ if }p=3;\\ \frac{3}{8}&\text{ if }p=5;\end{cases}$
@@
[…]
@@
Let $E_{A,B}\mathrel{\mathop{\ordinarycolon}}y^{2}=x^{3}+Ax+B$ be an elliptic curve such that either $\ell^{4}\nmid A$ or $\ell^{6}\nmid B$. Assume that choice of $\Sigma_{\ell}$ is made for every prime number. Then, we set $\Sigma$ to denote the tuple of conditions $(\Sigma_{\ell})_{\ell}$, where $\ell$ ranges over all prime numbers. We associate the family $\mathcal{F}_{\Sigma}$ of elliptic curves to the tuple $(\Sigma_{\ell})_{\ell}$, where $E_{A,B}\in\mathcal{F}_{\Sigma}$ if $(A,B)\in\Sigma_{\ell}$ for all $\ell$. The family $\mathcal{F}_{\Sigma}$ is then said to be defined by congruence conditions.
@@
$\operatorname{Inv}(\mathcal{F})\mathrel{\mathop{\ordinarycolon}}=\{(A,B)\mathrel{\mathop{\ordinarycolon}}E_{A,B}\in\mathcal{F}\}$
@@
Recall that the root number $\omega(E)$ of an elliptic curve $E_{/\mathbb{Q}}$ is the sign of the functional equation of the $L$-function $L(E,s)$ of $E$. It is widely believed that when the elliptic curves are ordered by height, the root numbers +1+1 and 1-1 occur equally often. The following result of Bhargava and Shankar [bhargavashankar15, Theorem 41] then gives the density of elliptic curves with rank 0.
@@
Let $p$ be a prime in $\{3,5\}$. Suppose $\mathcal{F}$ is a large family of elliptic curves such that exactly $50\%$ of the curves in $\mathcal{F}$, when ordered by height, have root number +1+1. Then if $p=3$ (resp. $p=5$) at least $25\%$ (resp. $37.5\%$) of the curves $E_{A,B}$ in $\mathcal{F}$, when ordered by height, have $\operatorname{Sel}_{p}(E_{A,B}/\mathbb{Q})=0$.
@@
The statement of the result above is a little stronger than that of [bhargavashankar15, Theorem 41] (resp. [bhargavashankar5, Proposition 40]) since it is asserted that $\operatorname{Sel}_{p}(E_{A,B}/\mathbb{Q})=0$. The proof of the mentioned Theorems rely on showing that this Selmer group is indeed 0 for at least $25\%$ (resp. $37.5\%$) of curves in $\mathcal{F}$. When $p=3$, we refer to p. 617, ll. 11-12 of [bhargavashankar15]. On the other hand, for $p=5$, see [bhargavashankar5, p.29, l. -7]. ∎
For an elliptic curve $E\mathrel{\mathop{\ordinarycolon}}=E_{A,B}\mathrel{\mathop{\ordinarycolon}}y^{2}=x^{3}+Ax+B$ with $A,B\in\mathbb{Z}$, denote by $E^{-1}\mathrel{\mathop{\ordinarycolon}}=E^{-1}_{A,B}\mathrel{\mathop{\ordinarycolon}}y^{2}=x^{3}+Ax-B$ the twist by 1-1 of $E_{A,B}$. Note that $\Delta(E^{-1})=\Delta(E)$ and $j(E^{-1})=j(E)$.
@@
[…]
There is a simple criterion for this at a prime $\ell\geq 3$ of multiplicative reduction. The elliptic curve $E=E_{A,B}$ has split multiplicative reduction precisely when $\left(\frac{6B}{\ell}\right)=1$ (see [Wong, Lemma 10]). One has that $\left(\frac{-1}{\ell}\right)=1$ if and only if $\ell\equiv 1\pmod{4}$. Therefore, we deduce that for any prime $\ell\geq 3$ at which $E$ has multiplicative reduction,
@@
    Let $\Delta^{\prime}(E)$ be the positive prime to $2$ part of $\Delta(E)$ (i.e., $\Delta^{\prime}(E)\mathrel{\mathop{\ordinarycolon}}=\frac{|\Delta(E)|}{2^{v_{2}(\Delta(E))}}$). Then, $\Delta^{\prime}(E)$ is squarefree and $\Delta^{\prime}(E)\equiv 1\pmod{4}$.
@@
Let $E$ be an elliptic curve in $\mathcal{E}^{0}$. Then, setting $K\mathrel{\mathop{\ordinarycolon}}$$\mathbb{Q}$, the Assumption 2.1 is satisfied for $\mathcal{E}^{0}$.
@@
  • Finally, let us consider $\ell=2$. By part (4) of Definition 3.12, $E$ has additive reduction at $2$ and $2\nmid j(E)$. It follows from [Wong, Lemma 12] that $\omega_{2}(E^{-1})=-\omega_{2}(E)$.
@@
Recall that the height of $E$, denoted $h(E)$, is defined to be $\operatorname{max}\{|A|^{3},B^{2}\}$. Thus, $h(E)$ remains unaltered under the twist by 1-1, i.e. $h(E)=h(E^{-1})$. Hence, Lemma 3.15 implies that exactly $50\%$ of the curves $E$ in $\mathcal{E}$ have root number +1+1, when ordered by height. We note that there is a natural short exact sequence
@@ 4.1. A suitable large family of elliptic curves
Clearly, the set $\mathcal{F}$ is a large family, and at the primes $T\mathrel{\mathop{\ordinarycolon}}=Z\cup\{2,p\}$, we have associated congruence conditions.
@@
for all $E\in\mathcal{F}$. That $E$ and $E^{-1}$ have additive reduction at $2$ follows from Tate’s algorithm (see [SilvermanAdvancedtopics, step 7, p. 367]). Since it is a standard procedure, we omit these details. The $j$-invariant is given as follows
@@
    Given $X>0$, we set $\mathcal{F}_{\ell}(X)\mathrel{\mathop{\ordinarycolon}}=\left\{(A,B)\in\mathcal{F}_{\ell}\mid\operatorname{max}\{|A|^{3},B^{2}\}\leq X\right\}$. We note that not all pairs $(A,B)\in\mathcal{F}_{\ell}$ are minimal, but the pairs $(A,B)\in\mathcal{F}$ are. We derive an asymptotic formula for $\mathcal{F}_{\ell}(X)$ as $X\rightarrow\infty$.
@@
    Next, we show that the asymptotic formulae for $\mathcal{F}_{\ell}(X)$ can be used to give an asymptotic formula for $\mathcal{F}(X)$. Using this formula, we derive an expression for the density of $\mathcal{F}$. Since $\mathcal{F}$ is contained in $\mathcal{E}$, this gives us lower bound for the lower density of $\mathcal{E}$.
In order to complete step (1) in an effective manner, we introduce the residue classes $\overline{\mathcal{F}}_{\ell}\subset(\mathbb{Z}/\ell^{n_{\ell}})^{2}$ associated to the local condition $\mathcal{F}_{\ell}$ for all primes. We then set $\delta_{\ell}\mathrel{\mathop{\ordinarycolon}}=\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}$. The asymptotic for $\mathcal{F}_{\ell}(X)$ is then given by
$\mathcal{F}_{\ell}(X)\sim\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}\times\#\{(A,B)\in\mathbb{Z}^{2}\mid\operatorname{max}\{|A|^{3},B^{2}\}\leq X\}\sim 4\delta_{\ell}X^{5/6}.$
@@
  • For $\ell\notin T$, we set $n_{\ell}=2$ and $\overline{\mathcal{F}}_{\ell}\subset(\mathbb{Z}/\ell^{2}\mathbb{Z})^{2}$ to consist of pairs $(\overline{A},\overline{B})$ for which $4\overline{A}^{3}+27\overline{B}^{2}\neq 0$.
@@
  • For $\ell=2$, we set $n_{\ell}=12$ and $\overline{\mathcal{F}}_{\ell}\subset(\mathbb{Z}/2^{12}\mathbb{Z})^{2}$ to consist of pairs $(\overline{A},\overline{B})$ where $(\overline{A},\overline{B})=(756,\pm 16)$.
@@
Let $\ell$ be a prime, we find that $\mathcal{F}_{\ell}(X)\sim 4\delta_{\ell}X^{5/6}$, where
[…]
@@
$\overline{\mathcal{F}}_{\ell}\subset(\mathbb{Z}/\ell^{n_{\ell}})^{2}$
@@
$\mathcal{F}_{\ell}(X)\sim\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}\right)\times 4X^{5/6}.$
We count $\overline{\mathcal{F}}_{\ell}$ in all cases.
  • First consider $\ell\notin T$. Recall that for $\ell\notin T$, $\mathcal{F}_{\ell}$ consists of all pairs $(A,B)\in\mathbb{Z}^{2}$ such that $\ell^{2}\nmid\Delta_{A,B}$. Thus, $\overline{\mathcal{F}}_{\ell}$ consists of all pairs $(\overline{A},\overline{B})\in(\mathbb{Z}/\ell^{2})^{2}$ such that $4\overline{A}^{3}\neq 27\overline{B}^{2}$. Note that $\ell\neq 2$ since $2\in T$. First assume that $\ell=3$ (this case potentially arises when $p=5$). In this case, the condition becomes $3\nmid\overline{A}$, and thus, $\overline{\mathcal{F}}_{\ell}=2\times 3^{3}$ and thus,
    $\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}\right)=\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{4}}\right)=\frac{2}{3}.$
    Next, we move on the case when $\ell\neq 3$. Set $(C,D)\mathrel{\mathop{\ordinarycolon}}=(-\overline{A}/3,\overline{B}/2)$ and count the number of pairs for which $C^{3}=D^{2}$. We consider two further cases. In the first case, $\ell\nmid C$, we set $(U,V)\mathrel{\mathop{\ordinarycolon}}=(C,D/U)$ and the equation becomes $U=V^{2}$. Thus, in this case, there are $(\ell^{2}-\ell)$ values of $V$ and $U$ is determined by $V$. On the other hand, if $\ell|C$ then $\ell|D$ as well and there are $\ell^{2}$ choices in this case. Thus, we get
    $\overline{\mathcal{F}}_{\ell}=\ell^{4}-(\ell^{2}-\ell)-\ell^{2}=\ell^{4}-2\ell^{2}+\ell.$
@@
    $\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{2n_{\ell}}}\right)=\left(\frac{\#\overline{\mathcal{F}}_{\ell}}{\ell^{4}}\right)=1-\frac{2}{\ell^{2}}+\frac{1}{\ell^{3}}.$